# SLIM: Sim-to-Real Legged Instructive Manipulation via Long-Horizon Visuomotor Learning

### v2

@@ Abstract
We present a low-cost legged mobile manipulation system that solves long-horizon real-world tasks, trained by reinforcement learning purely in simulation. This system is made possible by 1) a hierarchical design of a high-level policy for visual-mobile manipulation following instructions and a low-level policy for quadruped movement and limb control, 2) a progressive exploration and learning approach that leverages privileged task decomposition information to train the teacher policy for long-horizon tasks, which will guide an imitation-based student policy for efficient training of the high-level visuomotor policy, and 3) a suite of techniques for minimizing sim-to-real gaps.
   In contrast to previous approaches that use high-end equipment, our system demonstrates effective performance with more accessible hardware – specifically, a Unitree Go1 quadruped, a WidowX250S arm, and a single wrist-mounted RGB camera – despite the increased challenges of sim-to-real transfer. When fully trained in simulation, a single policy autonomously solves long-horizon tasks such as search, move, grasp, and drop-into, achieving nearly 80% success. This performance is comparable to that of expert human teleoperation on the same tasks but operates in a more efficient way, at 1.5 times the speed of human expert. The sim-to-real transfer is fluid across diverse indoor and outdoor scenes under varying lighting conditions. Finally, we discuss the key techniques that enable the entire pipeline, including efficient RL training and sim-to-real, to work effectively for legged mobile manipulation, and present their ablation results.
Figure 1: SLIM in Real. Snapshots of the same SLIM policy deployed in diverse real-world scenes, featuring significant variations in ground types, backgrounds, distractors, and other environmental factors. These scenes are not available in simulation training. The annotations at the bottom are added for understanding task progress and are not part of the input to the system.
@@ I Introduction
Until recently, robots have primarily been limited to static, highly repetitive tasks, often in industrial manufacturing. As the demand for autonomy grows, the need emerges for robots to navigate and interact with their environments - commonly referred to as mobile manipulators. Such systems can typically handle primitive tasks, such as searching, grasping, and transporting objects, to achieve overarching long-horizon goals.
There are several promising methodologies for training autonomous mobile manipulators: imitation learning from human demonstrations, reinforcement learning (RL) in the real world, and RL in simulation with sim-to-real transfer. Each of them has its pros and cons.
Imitation learning (IL) benefits from learning directly from human experts teleoperating the hardware, and thus reasonable behaviors and some success in a restricted domain can be achieved even with a handful of trajectories. However, diverse behavior data is required to improve the policy’s robustness even for a single task. This makes IL costly, especially if we further want to cover all the combinatorial variations of scenes and tasks for robot ability scaling.
Reinforcement learning (RL) can potentially generate an infinite amount of training data by allowing the robot to explore and interact with the environment. Furthermore, the resulting policy is often more optimal than a human teleoperating policy. However, directly training an RL policy in the real world can be challenging due to hardware maintenance, reward specification, and safety issues. Data scale is also largely limited by the difficulty of deploying a robot fleet for collecting data in parallel. As a result, mastering long-horizon tasks in the real world still poses a significant challenge.
RL in simulation could address the scaling, generalization, and safety issues by running massively parallel simulation with randomized scenes and tasks. But to successfully deploy the trained policy on a real robot, the challenge lies in sim-to-real gaps. Therefore, we attempt to develop effective and general techniques to bridge the gaps. This work focuses on training an end-to-end RL policy in simulation, and zero-shot deploying in the real world on a relatively low-cost quadruped based robot to solve long-horizon, intricate manipulation tasks. The complexity of accurate manipulation on legs, the variations of the physical environment, the scene and task setups, and the long task horizon are some of the factors that compound and increase the difficulty of RL training. On top of these factors, limited hardware reliability further exacerbates the sim-to-real gap, making this a daunting task.
We train an end-to-end model that is able to achieve long-horizon tasks consisting of searching, grasping, transporting and dropping primitive tasks (Figure 1). For robot mobility, we choose a quadruped instead of a wheeled base, as it has the potential to adapt to many different types of terrains. The choice of a quadruped base results in a large 19 DOF (12 leg joints, 6 arm joints, and 1 gripper) system, exacerbating the already difficult long-horizon RL exploration issue. To mitigate this, we utilize a hierarchical teacher-student design, consisting of a high-level mobile manipulation policy and low-level quadruped controller, and also a teacher-student training framework [6] for the high-level policy. To the best of our knowledge, this is the first demonstration of a low-cost real-world quadruped manipulator fully trained with RL in simulation, for solving long-horizon tasks.
The teacher policy has access to privileged information (e.g., structured and accurate object features) that is unavailable in the real world. To tackle the difficulties of long-horizon task learning, we also incorporate task decomposition as a form of privileged information, where boundaries between subtasks are defined. The student policy is then trained to distill the teacher policy while maximizing task rewards, but with access to only standard sensor data such as camera feed and proprioceptive state. We also condition the student visuomotor policy on language instructions, providing a natural way to the user for specifying the task to be executed by the robot.
We name our robotic system SLIM (Sim-to-Real Legged Instructive Manipulation). We reiterate that SLIM is a complete robotic system with fluid sim-to-real transfer. Unlike prior work [30, 53, 59], SLIM learns all modules in simulation from scratch, without the need for any pretrained model or real-world finetuning. SLIM relies on pure RL, and thus it does not require a single demonstration trajectory. Our contributions are as follows:
@@
    We develop a low-cost quadruped manipulation system SLIM. To our best knowledge, SLIM is the first end-to-end system for solving long-horizon real-world mobile manipulation tasks from RL alone.
@@
    To address the long-horizon RL exploration challenge, we propose a hierarchical framework for training SLIM’s high-level mobile manipulation policy and low-level quadruped controller, and adopt a teacher-student setup for high-level training efficiency.
@@
    We identify several techniques that are crucial for SLIM to achieve successful sim-to-real transfer with inexpensive hardware, as shown in real-world ablation studies.
  4. 4.
    We conduct an extensive set of experiments with real-world evaluations, achieving higher success rate and faster task completion than baselines. We also demonstrate robust task success across a wide range of indoor and outdoor environments in real-world.
@@ II-A Learning-based Quadruped Locomotion
More recently, there has been explosive progress in using learning-based approaches for achieving quadruped locomotion. In particular, sim-to-real reinforcement learning (RL) has arisen as a robust solution showcasing impressive feats such as parkour [61, 8, 4] and speeds of up to 3.9m/s [35]. Other works have also focused on exploiting the strengths of legged locomotion over their wheeled counterparts through utilizing the strong inherent coupling of proprioception with egocentric vision [16, 1] and active estimation of the environment [37, 28, 31]. Furthermore, minimal tuning is necessary compared to classical approaches [14] with massively parallelized simulations having been shown capable of learning gaits in just minutes [43]. Finally, data-driven legged locomotion has shown great promise in its ability to zero-shot generalize to new morphologies [13], surfaces [36, 28, 31] and agile skills [50, 27].
@@ II-B Legged Mobile Manipulation
Significant advancements in robust locomotion have further enabled researchers to push the boundaries of legged mobile manipulation. Building upon data-driven approaches, quadrupeds have been demonstrated pushing objects with their body [26], dribble around balls [27], and manipulate objects using egocentric [49] or calve mounted grippers [29]. Still, the most popular setup has been the traditional top-mounted 6DOF manipulator design [15, 18, 30, 59, 53, 34, 38, 45]. In addition to significant workplace expansion, manipulators offer other nuanced benefits, such as assisting with balance [22] and serving as an intuitive interface for collecting human demonstrations [18]. Classical approaches to solving top-mounted legged mobile manipulation consist of designing wholebody controllers [45] or even combining locomotion learning with model-based manipulation control [34]. Others simplify the need for low-level controllers by using high-level API calls ($\dot{x},\dot{\theta}$) provided by their respective quadruped robots to achieve graspability-aware policies [59], navigational pick-and-place [53], and language-conditioned mobile manipulation [40]. Complex tasks have also been achieved by leveraging teacher-student training setups in simulation for grasping [30] and door opening [60]. Finally, to expand the robot’s workspace even further, wholebody loco-manipulation has arisen where leg joints are actuated in a way to assist manipulation [15, 30, 38].
Most similar to our work, Liu et al. [30] introduced VBC, which trains a sim-to-real visuomotor whole-body loco-manipulation policy using a high-low hierarchical model and teacher-student training setup for the object pickup task. They achieve impressive results across a diverse range of objects by leveraging whole-body manipulation, though this approach comes with several key caveats. First, VBC requires the user to manually click on the target object to perform an initial segmentation using a third-party general-purpose model. This segmentation is then used to track the object through another third-party model. Consequently, to ensure consistent segmentation, the object must always remain within the frame. This requirement inherently limits the feasible starting state distribution of VBC, as the target object must be placed within a short distance in front of the robot and in view. Finally, due to the close proximity and visibility of the object, the VBC task horizon is relatively short, ending as soon as the object is lifted. In comparison,
@@
    SLIM is a complete and self-contained system. All modules–vision, high-level visual loco-manipulation, low-level motor control–are trained within a single framework, resulting in both minimal latency and full autonomy.
@@
    SLIM is fluid and intuitive. Conditioned on language, SLIM is able to rapidly accomplish tasks in a wide variety of scenes, requiring just a user language command.
@@
    Finally, SLIM is practical. With SLIM, we demonstrate impressive real-world, long-horizon task completion for a multi-stage mobile search and manipulation task.
@@ III Architecture Overview
Figure 2: Hierarchical Framework and Visuomotor Policy Pipeline (Left). Given a language instruction and various sensor information, the policy generates two control signals: 1) arm joint position targets and 2) a locomotion command. The first control signal is directly passed to the arm driver (passthrough), while the second signal is passed into a low-level locomotion policy to generate leg joint position targets. Low-Level and High-Level Teacher-Student Policy Training (Right). SLIM’s training pipeline can be decomposed into the following three sequential stages of policy training. First, a low-level locomotion policy is trained to track a specified locomotion command via RL. Next, a high-level teacher policy is trained with privileged information via RL to solve the long-horizon task. Finally, the student visuomotor policy is trained by distilling the teacher policy while maximizing task rewards, conditioned on visual inputs. Both the teacher and student policies use the same frozen locomotion policy from the first stage to execute their locomotion commands.
@@ III-A Observations and Actions
    A temporal stack of RGB images from a single wrist-mounted camera with a shape of $(N,H,W,3)$, where $N$ is the stack size, and $H$ and $W$ are the image height and width, respectively. The environment is always partially observed by the robot.
@@
    A temporal stack of proprioceptive state vectors of the robot with a shape of $(N,D)$, where $D$ is the dimensionality of the concatenation of proprioceptive reading of all joints. $\mathbf{s}_{\textrm{r}}\!=\![\mathbf{s}_{\textrm{leg}},\mathbf{s}_{\textrm{arm}}]$, where $\mathbf{s}_{\textrm{leg}}$ denotes the leg-related proprioceptive states and $s_{\textrm{arm}}$ the arm-related proprioceptive states. When used as the input to the low-level locomotion policy, we overload $\mathbf{s}_{\textrm{leg}}$ to denote the leg-related proprioceptive states stacked according to the low-level frame-stacking size.
@@
    A group of temporal object feature stacks, where each stack has a shape of $(N,Q_{m})$ and $0\leq m<M$ denotes the feature index. $M$ is the maximal number of task related objects in the scene. Note that object features are also partially observed and their visibility is always determined by the robot’s camera field of view.
@@
    A vector in $[-z,z]^{6}$, indicating the delta changes to the arm joint positions. $z$ is the maximally allowed joint position change for a control interval.
@@
    The target quadruped joint positions in $\mathbb{R}^{12}$.
In total, a complete action $\mathbf{a}$ has 19 dimensions.
@@ III-B Hierarchical Policy Structure
The system as a whole takes visual and instructional input and controls all joints of the quadruped and the arm, which is a daunting training task with high-dimensional input and output. We use a two-level hierarchical policy to divide the complexity of visual mobile manipulation training from legged locomotion training, similar to [30], illustrated in the left half of Fig. 2.
The high level takes visual, instructional, and proprioceptive inputs, and outputs a simple locomotion speed command and arm control command, while the low level takes proprioceptive input and outputs leg joint commands to track the locomotion command. The assumption is that quadruped locomotion control is largely independent of the high-level task semantics.
Intermediary Action.
To enable this two-level policy structure, we introduce an intermediary action as follows,
@@
$(\mathbf{l},\mathbf{o},\mathbf{s_{\textrm{r}}})\rightarrow\mathbf{a}_{\text{hi}}\triangleq(\mathbf{a_{\textrm{arm}}},g,\mathbf{c})\in\mathbb{R}^{9},$
@@
$(\mathbf{c},\mathbf{s_{\textrm{r}}})\rightarrow\mathbf{a_{\textrm{leg}}}\in\mathbb{R}^{12}.$
@@
Note that with this decomposition, the low-level policy no longer observes task specific inputs $(\mathbf{l},\mathbf{o})$, thus does not need to handle high dimensional image and language inputs.
Low-Level.
@@
The low-level policy is a quadruped controller that generates joint position targets for PD control to follow a task-agnostic 2D command $\mathbf{c}$ [36]. The low-level policy is trained in simulation using PPO [44] and transferred to real. An illustration of the training can be found in the first block in the right column of Fig. 2. It is worth noting that we randomly sample both command $\mathbf{c}$ and arm joint actions $\mathbf{a_{\textrm{arm}}}$ during low-level training. Since the arm mounted on the quadruped can have varying poses during the task, the low-level command following ability must generalize to various arm configurations. That is, the quadruped base has to learn to keep balance while achieving locomotion commands regardless of the arm’s current joint positions and movements. Similar to most sim2real approaches, we also randomize the simulation environment during training. Besides widely-used domain randomization parameters as in [15], we further randomize the simulated delay of each sensor and foot softness of the robot, to better adapt to variations in real deployment. Rewards design and full details of the training setup can be found in Appendix A-B.
High-Level.
@@
The high-level policy is responsible for outputting the command $\mathbf{c}$ and the arm control $(\mathbf{a_{\textrm{arm}}},g)$, conditioned on language $\mathbf{l}$ and the current stack of RGB images $\mathbf{o}$. As demonstrated in Fig. 2, the command $\mathbf{c}$ is forwarded to the low-level policy which then follows the received command for a number of time steps, before the high-level policy outputs the next command. For the manipulator, we opt to operate in joint space (as opposed to task space) and thus no inverse kinematics is needed. Once the low-level policy is trained, it is frozen and used as a base controller by the high-level policy. In the remainder of this paper, we will only talk about the high-level policy unless otherwise stated.
III-C Teacher-Student Training

### v3

@@ Abstract
We present a low-cost legged mobile manipulation system that solves long-horizon real-world tasks, trained by reinforcement learning purely in simulation. This system is made possible by 1) a hierarchical design of a high-level policy for visual-mobile manipulation following task instructions, and a low-level quadruped locomotion policy, 2) a teacher and student training pipeline for the high level, which trains a teacher to tackle long-horizon tasks using privileged task decomposition and target object information, and further train a student for visual-mobile manipulation via RL guided by the teacher’s behavior, and 3) a suite of techniques for minimizing the sim-to-real gap.
   In contrast to previous works that use high-end equipment, our system demonstrates effective performance with more accessible hardware – specifically, a Unitree Go1 quadruped, a WidowX-250S arm, and a single wrist-mounted RGB camera – despite the increased challenges of sim-to-real transfer. Trained fully in simulation, a single policy autonomously solves long-horizon tasks involving search, move to, grasp, transport, and drop into, achieving nearly 80% real world success. This performance is comparable to that of expert human teleoperation on the same tasks but significantly more efficient, operating at about 1.5$\times$ the speed. Finally, we perform extensive ablations on key techniques for efficient RL training and effective sim-to-real transfer, and demonstrate effective deployment across diverse indoor and outdoor scenes under various lighting conditions.
Figure 1: SLIM in Real. Snapshots of the same SLIM policy deployed in diverse real-world scenes, featuring significant variations in terrain, background, distractors, and other environmental factors. These scenes are not available in simulation training. The subtask annotations at the bottom are added for understanding task progress and are not part of the input to the system.
@@ I Introduction
Legged mobile manipulation combines a robotic manipulator with a legged mobile platform, enabling robots to perform a wide variety of complex tasks in diverse environments [16, 59, 41, 33, 43, 65]. Unlike stationary or wheeled manipulators, legged systems can adapt to uneven terrains, significantly expanding the scope of potential applications. Recent advancements in hardware and algorithms [16, 46, 39] have also made legged manipulators more accessible and cost-effective, encouraging their adoption in research and industry.
Despite its promise, legged mobile manipulation presents unique challenges. The expanded scope of the tasks, the long task horizon, the diverse scenes and terrains, and unstable legged base all compound the difficulty of achieving reliable performance.
One popular approach to tackle mobile manipulation is Imitation Learning (IL) from expert demonstrations [5, 18, 54, 4]. It is straightforward, leveraging human expertise to produce reasonable behaviors quickly with supervised training. However, IL relies heavily on large and comprehensive datasets to ensure generalization and robustness, making data collection resource-intensive, both in terms of human labor and hardware wear-and-tear. Moreover, its performance is upper-bounded by the quality of human demonstrations. These limitations become even more pronounced when addressing the complexity of legged manipulation.
Another paradigm in robotics learning is sim-to-real Reinforcement Learning (RL). This method bypasses the need for extensive real-world demonstration data by leveraging virtually infinite data available from simulation, ensuring robust policy training through diverse and comprehensive data [50, 26, 28, 46, 33, 11]. Furthermore, simulation allows safe exploration without physical wear-and-tear and is not constrained by the performance of demonstrations, unlike IL. However, the success of this paradigm hinges on tackling two key challenges: (1) effectively training RL policies in simulation over long horizons, and (2) addressing sim-to-real gaps, including those both in dynamics and vision. These challenges are further exacerbated by the additional complexities introduced by legged manipulators.
Based on sim-to-real RL, we present SLIM (Sim-to-Real Legged Instructive Manipulation), a standalone pipeline for training robotic policies entirely in simulation and deploying them zero-shot in the real world. SLIM employs a hierarchical policy structure that decouples task-dependent high-level visuomotor commands from low-level control (Section III-B). In high-level policy training, we utilize a teacher-student learning framework to enhance training efficiency, where a teacher policy leverages privileged information to guide the student. We train the teacher policy to solve long-horizon tasks via task decomposition and progressive learning (Section IV). We then train a student policy, conditioned on language and visual inputs, in RL guided by the teacher’s behavior (Section V). Together with a suite of carefully designed sim-to-real techniques (Section VII), including visual and dynamics randomization and low-level controller tuning, SLIM produces visuomotor policies that can maintain sim performance when transferred to real.
In this work, we implement SLIM on a Unitree Go1 quadruped with a top-mounted WidowX-250S manipulator and a wrist-mounted Intel RealSense D435 camera (only using the RGB stream), resulting in a large 19 degree-of-freedom (DOF) system. We focus on the long-horizon task consisting of multi-stage search, move to, grasp, transport, and drop into, achieving nearly 80% real-world success (Section VIII).
Our key contributions are as follows:
@@
    We develop a low-cost legged manipulation system SLIM. To the best of our knowledge, SLIM is the first end-to-end robotic system for solving long-horizon real-world legged manipulation tasks from sim-to-real RL alone.
@@
    We conduct an extensive set of real-world experiments with 400 episodes achieving human expert-level success ($\sim$80%) with $\sim$1.5$\times$ the speed of human expert teleoperation. We also demonstrate robust task success across a wide range of indoor and outdoor scenarios.
@@
    We identify crucial techniques for achieving successful sim-to-real transfer with inexpensive hardware, validated through extensive real-world ablations.
We plan to open source the code to facilitate future efforts on related research.
@@ II-A Learning-based Quadruped Locomotion
More recently, there has been explosive progress in using learning-based approaches for achieving quadruped locomotion. In particular, sim-to-real reinforcement learning (RL) has arisen as a robust solution showcasing impressive feats such as parkour [67, 8, 6] and high running speeds [38]. Other works have also focused on exploiting the strengths of legged locomotion over their wheeled counterparts through utilizing the strong inherent coupling of proprioception with egocentric vision [17, 1] and active estimation of the environment [40, 31, 34]. Furthermore, necessary tuning can be minimal compared to classical approaches [15], for instance, massively parallelized simulation has been shown capable of learning gaits in just minutes [46]. Finally, data-driven legged locomotion has shown great promise in its ability to zero-shot generalize to new morphologies [14], surfaces [39, 31, 34] and agile skills [56, 30].
@@ II-B Legged Mobile Manipulation
Significant advancements in robust locomotion have further enabled researchers to push the boundaries of legged mobile manipulation. Building upon data-driven approaches, quadrupeds have been demonstrated pushing objects with their body [29], dribbling around balls [30], and manipulating objects using egocentric [55] or calve mounted grippers [32]. Still, the most popular setup has been the traditional top-mounted 6DOF manipulator design [16, 20, 33, 65, 59, 37, 41, 48]. In addition to significant workplace expansion, manipulators offer other nuanced benefits, such as assisting with balance [24] and serving as an intuitive interface for collecting human demonstrations [20]. Classical approaches to solving top-mounted legged mobile manipulation consist of designing wholebody controllers [48] or even combining locomotion learning with model-based manipulation control [37]. Others simplify the need for low-level controllers by using high-level API calls provided by their respective quadruped robots to achieve graspability-aware policies [65], navigational pick-and-place [59], and language-conditioned mobile manipulation [43]. Complex tasks have also been achieved by leveraging teacher-student training setups in simulation for grasping [33] and door opening [66]. Finally, to expand the robot’s workspace even further, wholebody loco-manipulation has arisen where leg joints are actuated in a way to assist manipulation [16, 33, 41].
Most closely related to our work, Liu et al. [33] introduces a visual whole-body control approach (VBC), which trains a sim-to-real visuomotor whole-body loco-manipulation policy using a high-low hierarchical model and teacher-student training setup for the object pickup task, similar to our design. The system can grasp an impressive set of diverse objects by leveraging whole-body manipulation with capable hardware. VBC takes a few shortcuts to simplify the task. VBC requires the user to manually click on the target object to perform initial segmentation using a third-party vision model. This segmentation is then used to track the object through a third-party tracking model, which requires the object to stay within the camera view the whole time. Additionally, VBC uses a scripted policy to put the object into the basket on the back of the quadruped, once the object is lifted, making the task scope more limited and the task horizon much shorter. In comparison, SLIM offers the follow advantages:
@@
    SLIM is a complete and self-contained system. All modules–vision, high-level visual policy, low-level motor control–are trained in our simulation environment integrated within a single framework without any third party modules, resulting in a lean 19 million parameter neural network system, allowing both minimal latency and full autonomy.
@@
    SLIM is fluid and intuitive to use. Given a language instruction, SLIM is able to rapidly accomplish tasks in a wide variety of scenes.
@@
    Finally, SLIM is practical. SLIM can accomplish complex long-horizon real-world tasks autonomously from the beginning to end, as apposed to some of the existing methods that requires human involvement (e.g. manually clicking for segmentation) and can only automate part of the task with restrictions on the feasible starting state distribution [33]
@@ III Architecture Overview
Figure 2: Hierarchical Framework and Visuomotor Policy Pipeline (Left). Given language instruction and sensor inputs, the high-level policy generates a set of two control signals: 1) the arm and gripper control signals, and 2) the locomotion command. The arm control signals are directly passed to the arm driver, and the locomotion command is passed to the low-level policy to control the leg joints of the quadruped. Low-Level Training (Top Right), and High-Level Teacher and Student Training (Middle and Bottom Right). SLIM training is divided into three sequential stages. First, the low-level locomotion policy is trained via RL to follow a sampled linear and angular velocity command. Second, the high-level teacher policy is trained via RL with privileged low-dimensional state input to solve the long-horizon task. Finally, the student visuomotor policy is trained by distilling the teacher behavior while maximizing task rewards, using visual, sensory, and language instruction as input. Both the teacher and student policies command the same frozen low-level policy produced by the first stage. The teacher is only run in simulation and the student can be deployed in real.
Fig. 2 illustrates the hierarchical (high-low policy) framework (left) and the teacher-student training pipeline (right) used in SLIM. Given language instruction and sensor inputs, the high-level policy generates intermediate control signals that will be used as input to the low-level policy. A low-level locomotion policy is trained to track a specified locomotion command via RL and is frozen after training. Then a high-level teacher policy is trained with privileged information via RL to solve the long-horizon task, and a student visuomotor policy is trained by distilling the teacher policy while maximizing task rewards, conditioned on visual inputs and language.
@@ III-A Observations and Actions
    A temporal stack of RGB images from a single wrist-mounted camera with a shape of $(N,H,W,3)$, where $N$ is the stack size, and $H$ and $W$ are the image height and width, respectively. The environment is always partially observed by the robot, and most of the workspace will not be contained in this camera view during the task.
@@
    A temporal stack of proprioceptive state vectors of the robot with a shape of $(N,D)$, where $D$ is the dimensionality of the concatenation of proprioceptive readings of all joints. $\mathbf{s}_{\textrm{r}}\!=\![\mathbf{s}_{\textrm{leg}},\mathbf{s}_{\textrm{arm}}]$, where $\mathbf{s}_{\textrm{leg}}$ denotes the leg-related proprioceptive state (leg joint positions) and $s_{\textrm{arm}}$ the arm proprioceptive state (arm joint positions).
@@
    A group of temporal stacks of privileged object features (e.g., positions, scales, categories, etc.), where each stack has a shape of $(N,Q_{m})$ and $0\leq m<M$ denotes the object index. $M$ is the maximal number of task related objects in the scene. Note that object features are also partially observed and their visibility is always determined by the robot’s camera field of view. The features are set to zero if they are not in the view. This is to ensure that the object visibility is consistent between the guidance policy and the robot policy, and avoid any information gap for policy distillation.
@@
    A vector in $[-z,z]^{I}$, indicating the delta changes to the arm joint positions. $z$ is the maximally allowed joint position change for a control interval. $I$ is the number of actuated arm joints, excluding the gripper joint.
@@
    The target quadruped joint positions in $\mathbb{R}^{J}$. $J$ is the number of actuated leg joints.
In total, a complete action $\mathbf{a}$ has $(I+1+J)$ dimensions.
@@ III-B Hierarchical Policy Structure
The overall system takes visual and instructional input and controls all joints of the quadruped and the arm, which involves a non-trivial training task with high-dimensional input and output. We use a two-level hierarchical policy to divide the complexity of visual mobile manipulation training from legged locomotion training, illustrated in the left part of Fig. 2, similar to [33].
The high level policy takes visual, instructional, and proprioceptive inputs, and outputs a locomotion speed command and arm control command. The low level policy tracks the locomotion command from the high level, using proprioceptive inputs to command the quadruped’s leg joints. The assumption is that given the locomotion command, quadruped locomotion control is largely independent of the high-level task semantics.
Intermediate Action.
The following intermediate action separates the high and low-level policies,
@@
$(\mathbf{l},\mathbf{o},\mathbf{s_{\textrm{r}}})\rightarrow\mathbf{a}_{\text{hi}}\triangleq(\underbrace{[\mathbf{a_{\textrm{arm}}},g]}_{\textrm{arm and gripper control}},\underbrace{{\mathbf{{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}[}c{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}]}}}}_{\textrm{locomotion command}})\in\mathbb{R}^{I+3},$
@@
$(\mathbf{c},\mathbf{s_{\textrm{leg}}})\rightarrow\mathbf{a_{\textrm{leg}}}\in\mathbb{R}^{J}.$
@@
Note that with this decomposition, the low-level policy no longer observes task specific inputs $(\mathbf{l},\mathbf{o})$ and thus, does not need to process high-dimensional image and language inputs.
Low Level.
@@
The low-level policy is a quadruped controller that generates joint position targets for PD control to follow a task-agnostic 2D linear and angular-velocity command $\mathbf{c}$ [39] and is trained in simulation using PPO [47]. An illustration of the training can be found in the top row of the right column of Fig. 2. It is worth noting that during low-level training, the same embodiment as shown in Fig. 2 (arm mounted on top of the quadruped) is used in low-level training. We randomly sample both command $\mathbf{c}$ and arm joint actions $\mathbf{a_{\textrm{arm}}}$ during training. Since the arm mounted on the quadruped can have varying poses during the task, this enforces the low-level command following ability to generalize to various arm configurations. That is, the quadruped base has to learn to keep balance while achieving locomotion commands regardless of the arm’s current joint positions and movements. We intentionally remove arm joint state and joint command from low level input, see details in Appendix B-C. Similar to most sim-to-real approaches, we also randomize the simulation environment during training. Besides widely-used domain randomization parameters as in [16], we further randomize the simulated delay of each sensor and foot softness of the robot to better adapt to variations in real deployment. Full details of the training setup and reward design can be found in Appendix A-C.
High Level.
@@
The high-level policy is responsible for outputting the locomotion command $\mathbf{c}$ and the arm controls $(\mathbf{a_{\textrm{arm}}},g)$, given language instruction $\mathbf{l}$, proprioceptive state $\mathbf{s}_{\textrm{r}}$, and the current stack of RGB observations $\mathbf{o}$. As shown in Fig. 2, the command $\mathbf{c}$ is forwarded to the low-level policy, which then follows the received command for a number of time steps, before the high-level policy outputs the next command. For the manipulator, we opt to operate in joint space, as opposed to task space or end-effector pose, so we do not have to worry about inverse kinematics computation. Once the low-level policy is trained, it is frozen and used as a base controller by the high-level (teacher or student) policies. In the remainder of main text, we only talk about the high-level policies unless otherwise stated.
III-C Teacher-Student For High-Level Training